OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–2, –3) and is parallel to the line y = x + 9? x – y = 1 x + y = 1 x – y = –5 x + y = –5

OpenStudy (anonymous):

What do we know about two parallel lines? What do they have in common?

OpenStudy (anonymous):

they never cross eachj othe

OpenStudy (anonymous):

Right, but mathematically speaking why don't they cross?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

That's okay. They have the same slope. So if they have the same slope and they give us the equation to the second line, we automatically know the slope of our line, right?

OpenStudy (anonymous):

Our line being the first one they want us to find.

OpenStudy (anonymous):

So, y=mx+b where m is the slope. What's the slope of both lines?

OpenStudy (anonymous):

oh okay so it would be slope for 9

OpenStudy (anonymous):

Not quite. Our given equation is y=1x+9, so our slope is 1, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

haha so then given slope of 1 and passing through the point (-2,-3). You have a basic equation given as y=1x+b. We need to find b, so let's substitute our point's x and y-values: \[(-3)=1(-2)+b\] then solve for b.

OpenStudy (anonymous):

how do solve it?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Sorry, was in class. Simple algebra from there, -3=-2+b, so b=-1. Now you know everything to write your equation in y=mx+b format.